OpenStudy (thefluffmuncher):

3. Solve by the linear combination method (with or without multiplication). 4g + 5h = 4 8g – 10h = 4 (Points : 1)

OpenStudy (thefluffmuncher):

i solved it but it gave me 0=4

OpenStudy (perl):

what is the linear combination method

OpenStudy (thefluffmuncher):

i their is not option that says it has no soultion am i doing something wrong

OpenStudy (thefluffmuncher):

solution*

OpenStudy (perl):

if you multiply the first equation by 2

OpenStudy (perl):

4g + 5h = 4 8g - 10h = 4 multiply first equation by 2 8g + 10h = 8 8g -10h = 4 add the two equations 16g = 12 g= 12/16= 3/4

OpenStudy (thefluffmuncher):

maybe it came out wrong because i multiplied it by a negative 2

OpenStudy (perl):

4g + 5h = 4 8g - 10h = 4 multiply first equation by -2 -8g -10h = -8 8g -10h = 4 add the two equations -20h = -4 h= 4/20= 1/5

OpenStudy (thefluffmuncher):

thanks for helping me :)